### v3

# Travelling waves for Maxwell’s equations in nonlinear and symmetric media

Abstract.
We look for travelling wave fields
$E(x,y,z,t)=U(x,y)\cos(kz+\omega t)+\widetilde{U}(x,y)\sin(kz+\omega t),\quad(x,y,z)\in\mathbb{R}^{3},\,t\in\mathbb{R},$
satisfying Maxwell’s equations in a nonlinear and cylindrically symmetric medium. We obtain a sequence of solutions with diverging energy consisting of transverse magnetic field modes. In addition, we consider a general nonlinearity, controlled by an N-function.
Key words and phrases:
Maxwell equations, Kerr nonlinearity, curl-curl problem, travelling waves, N-functions, variational methods
2000 Mathematics Subject Classification:
Primary: 35J20, 58E15; Secondary: 47J30, 35Q60
Introduction
We are interested in self-trapped beams of light propagating in a nonlinear dielectric and cylindrically symmetric medium by means of solutions to the system of Maxwell’s equations coupled with a nonlinear constitutive relation between the electric field and the electric displacement field. More precisely, we are looking for travelling electric fields of the form
(1.1) $E(x,y,z,t)=U(x,y)\cos(kz+\omega t)+\widetilde{U}(x,y)\sin(kz+\omega t),$
where $U,\widetilde{U}\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ are the profiles of the travelling waves, $\omega>0$ is the temporal frequency and $k\in\mathbb{R}\setminus\{0\}$ the spatial wave number in the direction of propagation.
In modelling optical nonlinear materials [18], it is natural to seek solutions having cylindrical symmetry, and in physics literature one considers usually transverse electric field modes (TE modes), or tranverse magnetic field modes (TM modes) [9]. If the electric field is perpendicular to the direction of propagation, i.e.
(1.2) $U(x,y)=\frac{\beta(r)}{r}\begin{pmatrix}-y\\ x\\ 0\end{pmatrix},\quad(x,y)\in\mathbb{R}^{2},\;r=|(x,y)|,$
then $E=U(x,y)\cos(kz+\omega t)$ is the cylindrically symmetric TE mode ($\widetilde{U}=0$), and the system of Maxwell’s equations can be reduced to the study of a nonlinear second order differential equation with boundary conditions at infinity, see Stuart [20, 21] and the references therein. If the magnetic field $B$ is transverse to the direction of propagation, and takes the form
(1.3) $B=\frac{\beta(r)}{r}\begin{pmatrix}-y\\ x\\ 0\end{pmatrix}\cos(kz+\omega t),$
then the equations governing these transverse magnetic field modes are more complicated than those for TE modes because of the form of the constitutive relation in a nonlinear optical medium. However, as observed in [24, 9], when studying TM-modes, the electric field takes the form (1.1) with two profiles $U$ and $\widetilde{U}$, with possibly $U\neq 0$ and $\widetilde{U}\neq 0$.
In order to find the exact propagation of the electromagnetic field in a nonlinear medium, we consider a nonlinear polarization of the form $\chi(\langle|E|^{2}\rangle)E,$ where $\chi$ is the scalar nonlinear susceptibility, which depends only on the time averaged intensity of $E$ over one period $T=2\pi/\omega$, that is, $\langle|E(x,y,z)|^{2}\rangle=\frac{1}{T}\int_{0}^{T}|E(x,y,z,t)|^{2}\,dt$, see [19]. According to Maxwell’s equations, one has to solve the nonlinear electromagnetic wave equation
(1.4) $\nabla\times\Big(\frac{1}{\mu}\nabla\times E\Big)+\partial_{tt}\left(\varepsilon E+\chi(\langle|E|^{2}\rangle)E\right)=0\quad\hbox{for }(x,y,z,t)\in\mathbb{R}^{3}\times\mathbb{R},$
where $\varepsilon$ is the permittivity of the medium, $\mu$ is the magnetic permeability. For simplicity, we assume that $\mu=1$; then, in absence of charges and currents, Faraday’s law $\nabla\times E+\partial_{t}B=0$ determines the magnetic induction $B$, and Ampére’s law $\nabla\times B=\partial_{t}D$ together with $\mathop{\mathrm{div}\,}D=0$ and $\mathop{\mathrm{div}\,}B=0$ shows that Maxwell’s equations are satisfied, where we consider the linear magnetic material law $B=\mu H=H$ and the nonlinear electric material law $D=\varepsilon(x,y)E+\chi(\langle|E|^{2}\rangle)E$, see [20, 21, 1, 14]. Altogether, we find the exact propagation of the electromagnetic field in a nonlinear medium according to Maxwell’s equations. Recall that in cylindrically symmetric media, $\varepsilon$, $U$, $\widetilde{U}$ depend only on $r=|(x,y)|$.
Problem (1.4) has been studied in cylindrically symmetric media in a series of previous works [22, 20, 21, 24, 23, 25, 26], where TE and TM modes have been obtained for asymptotically constant susceptibility $\chi$, i.e. in the so-called saturation effect. In this work, however, we want to treat the probably most common type of nonlinearity in the physics and engineering literature, i.e. the Kerr nonlinearity
$\chi(\langle|E^{2}|\rangle)E=\chi^{(3)}\langle|E|^{2}\rangle E,\quad\chi^{(3)}>0.$
Up to our knowledge, the first analytical work in this context is due to McLeod, Stuart and Troy [11], where TE modes were obtained. Namely, they showed that $\beta(r)$ must satisfy
(1.5) $\beta^{\prime\prime}+\frac{1}{r}\beta-\frac{1}{r^{2}}\beta+\beta^{3}-\beta=0,\quad\beta(0)=\beta(+\infty)=0$
(see also Pohl [16]), where $\beta$ is as in (1.3), and for a given integer $n\geq 1$, (1.5) has at least one solutions with precisely $n$ zeros in $(0,+\infty)$ whose derivative has precisely $n+1$ zeros. Then, up to some constant $\varepsilon$ and $\chi^{(3)}$, this leads to infinitely many different TE modes $E=U(r)\cos(kz+\omega t)$ with (1.2). Observe that a TE mode of this form is divergence free and $\widetilde{U}=0$.
The aim of this work is to find infinitely many fields $E$ of the form (1.1) which need not be divergence free, hence different from TE modes. Moreover, we admit a permittivity depending on $r$, $U$ need not be of the form (1.2), and the ODE methods do not seem to be applicable. Our approach is variational, and recall that the first author and Reichel in [14] have recently observed that the search for travelling waves of the form (1.1) satisfying (1.4) leads to the following nonlinear elliptic problem
(1.6) $L\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}-\omega^{2}\epsilon(x,y)\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}=\omega^{2}\chi\left(\frac{1}{2}(|U|^{2}+|\widetilde{U}|^{2})\right)\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}$,
where
$L=\begin{pmatrix}-\partial_{yy}+k^{2}&\partial_{xy}&0&0&0&k\partial_{x}\\ \partial_{xy}&-\partial_{xx}+k^{2}&0&0&0&k\partial_{y}\\ 0&0&-\partial_{xx}-\partial_{yy}&k\partial_{x}&k\partial_{y}&0\\ 0&0&-k\partial_{x}&-\partial_{yy}+k^{2}&\partial_{xy}&0\\ 0&0&-k\partial_{y}&\partial_{xy}&-\partial_{xx}+k^{2}&0\\ -k\partial_{x}&-k\partial_{y}&0&0&0&-\partial_{xx}-\partial_{yy}\end{pmatrix}$
is the second-order differential operator $L\colon\operatorname{dom}(L)\subset L^{2}(\mathbb{R}^{2})^{6}\to L^{2}(\mathbb{R}^{3})^{6}$, which is elliptic and self-adjoint, and
$\operatorname{dom}(L):=\left\{\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in L^{2}(\mathbb{R}^{2})^{6}:L\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in L^{2}(\mathbb{R}^{2})^{6}\right\},$
where $L((U,\widetilde{U})^{T})$ is defined in the sense of distributions. In view of [14, Theorem 2.3], the spectrum $\sigma(L)$ equals $\{0\}\cup[k^{2},\infty)$, where 0 is an eigenvalue of infinite multiplicity and $k^{2},\infty)$ consists of absolutely continuous spectrum. In [14], periodic media have been considered, but we borrow some ideas of the functional setting from this work to investigate the cylindrically symmetric situation.
We look for solutions $u\colon\mathbb{R}^{2}\to\mathbb{R}^{6}$ to the following slight generalization of (1.6) given by
(1.7) $Lu-V(x)u=f(u)\quad\hbox{for }x\in\mathbb{R}^{2}$
involving the operator $L$, where we assume $f(u)=F^{\prime}(u)$. From now on, the spacial variable in $\mathbb{R}^{2}$ will be denoted by $x$ instead of $(x,y)$. Observe that, if
(1.8) $V(x)=\omega^{2}\varepsilon(x),\quad F(u)=\omega^{2}\chi\Big(\frac{1}{2}|u|^{2}\Big),\quad u\in\mathbb{R}^{6},$
then (1.7) leads to (1.6) and we obtain the exact propagation of the travelling electromagnetic waves $(E,B)$, where $E$ is given by (1.1) and $B$ is provided by Faraday’s law and a subsequent time integration.
By a (weak) solution to (1.7) we mean a critical point of the functional
(1.9) $J(u):=\frac{1}{2}b_{L}(u,u)-\frac{1}{2}\int_{\mathbb{R}^{2}}V(x)|u|^{2}\,dx-\int_{\mathbb{R}^{2}}F(u)\,dx$
defined on a Banach space $X\subset L^{2}(\mathbb{R}^{2})^{6}$ given later, where $p>2$. Here, $b_{L}(\cdot,\cdot)$ is the bilinear form associated with $L$, i.e. $b_{L}(u,\varphi)=\int_{\mathbb{R}^{2}}\langle Lu,\varphi\rangle\,dx$ for all $u\in\operatorname{dom}(L)$ and all $\varphi\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$.
Now, for $\alpha,\widetilde{\alpha}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})$ and for $\beta,\widetilde{\beta}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{3}$ let us denote
$\begin{array}[]{ll}\overset{\circ}{\nabla}\begin{pmatrix}\alpha\\ \widetilde{\alpha}\end{pmatrix}:=\begin{pmatrix}\partial_{x}\alpha\\ \partial_{y}\alpha\\ k\widetilde{\alpha}\\ \partial_{x}\widetilde{\alpha}\\ \partial_{y}\widetilde{\alpha}\\ -k\alpha\end{pmatrix}\colon\mathbb{R}^{2}\to\mathbb{R}^{6},\qquad\overset{\circ}{\nabla}\times\begin{pmatrix}\beta\\ \widetilde{\beta}\end{pmatrix}:=\begin{pmatrix}\partial_{y}\beta_{3}-k\widetilde{\beta}_{2}\\ k\widetilde{\beta}_{1}-\partial_{x}\beta_{3}\\ \partial_{x}\beta_{2}-\partial_{y}\beta_{1}\\ \partial_{y}\widetilde{\beta}_{3}+k\beta_{2}\\ -k\beta_{1}-\partial_{x}\widetilde{\beta}_{3}\\ \partial_{x}\widetilde{\beta}_{2}-\partial_{y}\widetilde{\beta}_{1}\end{pmatrix}\colon\mathbb{R}^{2}\to\mathbb{R}^{6},\end{array}$
and observe that $L=\overset{\circ}{\nabla}\times\overset{\circ}{\nabla}\times$, $\overset{\circ}{\nabla}\times\overset{\circ}{\nabla}=0$, and thus $w:=\overset{\circ}{\nabla}\begin{pmatrix}\alpha\\ \widetilde{\alpha}\end{pmatrix}\in\ker(L)$. Moreover, $J$ may be unbounded above and below and its critical points may have infinite Morse index due to the infinite-dimensional kernel of $L$.
Due to the strongly indefinite nature of the problem and in order to control a large class of nonlinear phenomena, we need to introduce the notion of $N$-functions. We assume that there exists an N-function $\Phi\colon\mathbb{R}\to[0,\infty)$ with the following properties.
  • (N1)
    $\Phi$ satisfies the $\Delta_{2}$ and $\nabla_{2}$ conditions globally [^1] Such conditions are defined in Section 2..
  • (N2)
    $\lim_{t\to 0}\frac{\Phi(t)}{t^{2}}=0$.
  • (N3)
    $\lim_{t\to\infty}\frac{\Phi(t)}{t^{2}}=\infty$.
Let us collect assumptions about $V$ and $F$:
  • (V)
    $V\in L^{\infty}(\mathbb{R}^{2})$ is radially symmetric, positive, bounded away from 0, and $\operatorname*{ess\,sup}V<k^{2}$.
  • (F0)
    $F\colon\mathbb{R}^{6}\to\mathbb{R}$ is convex, nonnegative, and of class ${\mathcal{C}}^{1}$.
  • (F1)
    $f(u)=o(|u|)$ as $u\to 0$.
  • (F2)
    There exist $c_{1}>0$ such that for every $u\in\mathbb{R}^{6}$
    $|f(u)|\leq c_{1}\bigl(1+\Phi^{\prime}(|u|)\bigr)\quad\text{for all }u\in\mathbb{R}^{6}.$
  • (F3)
    There exists $\gamma>2$ and $c_{2}>0$ such that for every $u\in\mathbb{R}^{6}$
    $\frac{1}{\gamma}\langle f(u),u\rangle\geq F(u)\geq c_{2}\Phi(|u|).$
Here and in the sequel, $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product in $\mathbb{R}^{N}$, $N\geq 1$. Our model example is $F(u)=\frac{1}{p}|u|^{p}$ with $p>2$, where $p=4$ is responsible for the Kerr effect, and then the $N$-function is of power type $|u|^{p}$. In general, we can admit more examples, e.g.
(1.10) $f(u)=\begin{cases}|u|^{q-2}u\ln(1+|u|),\quad\hbox{ for }|u|>1,\\ |u|^{p-2}u\ln(2),\quad\hbox{ for }|u|\leq 1,\end{cases}$
where $p,q>2$. Then, $\Phi(t)=F(t)=\int_{0}^{t}f(s)\,ds$, (F1)–(F3) are satisfied, but $F$ cannot be controlled from both above and below by any power-type $N$-function.
Let $g\in{\mathcal{S}}{\mathcal{O}}(2)$ and let $u\colon\mathbb{R}^{2}\to\mathbb{R}^{6}$. We define
$\widetilde{g}:=\begin{pmatrix}g&0&0&0\\ 0&1&0&0\\ 0&0&g&0\\ 0&0&0&1\end{pmatrix}\in\mathbb{R}^{6\times 6}$
and
$(g\star u)(x):=\widetilde{g}u(g^{-1}x)$
for a.e. $x\in\mathbb{R}^{2}$.
Theorem 1.1.
Suppose that (V), (F0)–(F3) are satisfied and $F$ is radial. Then there exist infinitely many solutions to (1.7) of the form $u_{n}=v_{n}+w_{n}$ such that $J(u_{n})\to\infty$ as $n\to\infty$, $v_{n}\in H^{1}(\mathbb{R}^{2})^{6}$, $v_{n}\neq 0$, $w_{n}\in L^{2}(\mathbb{R}^{2})^{6}$, $\Phi(w_{n})\in L^{1}(\mathbb{R}^{2})$, $Lw_{n}=0$, and $g\star u_{n}=u_{n}$. Moreover, each $u_{n}=\begin{pmatrix}U_{n}\\ \widetilde{U}_{n}\end{pmatrix}$ has the profiles of the form
(1.11) $U_{n}=\frac{\alpha_{n}(x)}{|x|}\begin{pmatrix}x_{1}\\ x_{2}\\ 0\end{pmatrix}+\gamma_{n}(x)\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\quad\widetilde{U}_{n}=\frac{\widetilde{\alpha}_{n}(x)}{|x|}\begin{pmatrix}x_{1}\\ x_{2}\\ 0\end{pmatrix}+\widetilde{\gamma}_{n}(x)\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$,
for some radial functions $\alpha_{n},\widetilde{\alpha}_{n},\gamma_{n},\widetilde{\gamma}_{n}:\mathbb{R}^{2}\to\mathbb{R}$.
Observe that the solutions obtained in Theorem 1.1, with the first profiles of the form (1.11), cannot be TE-modes, i.e. $U$ of the form (1.2) and $\widetilde{U}=0$. Therefore, we obtain a sequence of solutions with diverging energy that is different from that obtained by McLeod, Stuart, and Troy in [11]. Moreover, if the electric field $E$ is of the form (1.1) with profiles (1.11), then the magnetic induction
$B=\nabla\times E=\begin{pmatrix}\partial_{x_{2}}\gamma_{n}-\widetilde{\alpha}_{n}x_{2}k\\ -\partial_{x_{1}}\gamma_{n}+\widetilde{\alpha}_{n}x_{1}k\\ 0\end{pmatrix}\cos(kx_{3}+\omega t)+\begin{pmatrix}\alpha_{n}x_{2}k+\partial_{x_{2}}\widetilde{\gamma}_{n}\\ -\alpha_{n}x_{1}k-\partial_{x_{1}}\widetilde{\gamma}_{n}\\ 0\end{pmatrix}\sin(kx_{3}+\omega t)$
is transverse to the direction of propagation, hence each $u_{n}$ is a TM-mode. Moreover the total electromagnetic energy per unit interval on the $x_{3}$-axis is finite, i.e
(1.12) ${\mathcal{L}}(t):=\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{a}^{a+1}\langle E,D\rangle+\langle B,H\rangle\,dx_{3}\,d(x_{1},x_{2})<+\infty,$
which is important in the study of self-guided beams of light in nonlinear media; see e.g. [20, 21].
The paper is organized as follows. In Section 2 we build the functional setting for the problem and recall some properties of the operator $L$. Our variational approach based on Cerami sequences is developed in Section 3. In the last Section 4 we introduce symmetric group actions which guarantee the profiles of the form (1.11) and exclude TE-modes. We also prove that ${\mathcal{L}}(t)$ is finite.
2. Variational setting
We introduce the following notation. If $u=\begin{pmatrix}U\\ {\tilde{U}}\end{pmatrix}\in\mathbb{R}^{6}$, then $|u|:=\big(|U|^{2}+|{\tilde{U}}|^{2}\big)^{1/2}$. In the sequel, $\langle\cdot,\cdot\rangle_{2}$ denotes the inner product in $L^{2}(\mathbb{R}^{2})^{6}$ and $|\cdot|_{q}$ denotes the usual $L^{q}$-norm for $q\in[1,+\infty]$. Furthermore, we denote by $C$ a generic positive constant which may vary from one inequality to the next. We always assume that $k\neq 0$. By $\lesssim$ we denote the inequality $\leq$ up to a positive multiplicative constant.
Now, let us recall from [17] some basic facts about N-functions and Orlicz spaces. A function $\Phi\colon\mathbb{R}\to[0,\infty)$ is called an N-function if and only if it is even, convex, and satisfies
$\lim_{t\to 0}\frac{\Phi(t)}{t}=0,\quad\lim_{t\to\infty}\frac{\Phi(t)}{t}=\infty,\quad\text{and}\quad\Phi(t)>0\text{ for all }t>0.$
We can associate with it a second function $\Psi\colon\mathbb{R}\to[0,\infty)$ defined as
$\Psi(t):=\sup\left\{s|t|-\Phi(s):s\geq 0\right\},$
which is again an N-function and is called the complementary function of $\Phi$. It is easy to check that $\Phi$ is the complementary function of $\Psi$.
We say that $\Phi$ satisfies the $\Delta_{2}$ condition globally if and only if there exists $K>0$ such that
$\Phi(2t)\leq K\Phi(t)\quad\text{for all }t\in\mathbb{R}.$
We say it satisfies the $\nabla_{2}$ condition globally if and only if there exists $K^{\prime}>1$ such that
$\Phi(K^{\prime}t)\geq 2K^{\prime}\Phi(t)\quad\text{for all }t\in\mathbb{R}.$
Recalling that convex functions are differentiable almost everywhere, we give the following characterization for the $\Delta_{2}$ and $\nabla_{2}$ conditions (see [17, Theorem II.III.3]).
Lemma 2.1.
The following properties are equivalent.
  • $\Phi$ satisfies the $\Delta_{2}$ condition globally.
  • There exists $\kappa>1$ such that $t\Phi^{\prime}(t)\leq\kappa\Phi(t)$ for a.e. $t\in\mathbb{R}$.
  • There exists $\kappa^{\prime}>1$ such that $t\Psi^{\prime}(t)\geq\kappa^{\prime}\Psi(t)$ for a.e. $t\in\mathbb{R}$.
  • $\Psi$ satisfies the $\nabla_{2}$ condition globally.
The set
$L^{\Phi}:=\left\{u\colon\mathbb{R}^{2}\to\mathbb{R}^{6}:\Phi\circ|u|\in L^{1}(\mathbb{R}^{2})\right\}$
is a vector space if $\Phi$ satisfies the $\Delta_{2}$ condition globally. In this case, it is called an Orlicz space. It becomes a Banach space if endowed with the norm
$|u|_{\Phi}:=\inf\left\{\alpha>0:\int_{\mathbb{R}^{2}}\Phi\left(\frac{|u|}{\alpha}\right)\,dx\leq 1\right\},$
and it is reflexive if $\Phi$ satisfies the $\Delta_{2}$ and $\nabla_{2}$ conditions globally, with $L^{\Psi}$ begin its dual space (see [17, Theorem IV.I.10 and Corollary IV.II.9]). We can define likewise the space of scalar functions
$L^{\Phi}(\mathbb{R}^{2}):=\left\{u\colon\mathbb{R}^{2}\to\mathbb{R}:\Phi\circ u\in L^{1}(\mathbb{R}^{2})\right\}$
and the corresponding norm, still denoted $|\cdot|_{\Phi}$. Then, $L^{\Phi}=L^{\Phi}(\mathbb{R}^{2})^{6}$ and their norms are equivalent (cf. e.g. [15, Lemma 2.1]).
An important property is the following one.
Lemma 2.2.
If $\Phi$ satisfies the $\Delta_{2}$ condition globally, then there exists $C>0$ such that
$\Psi\bigl(\Phi^{\prime}(t)\bigr)\leq C\Phi(t)$
for every $t\in\mathbb{R}$. In particular, if $u\in L^{\Phi}$, then $\Phi^{\prime}\circ|u|\in L^{\Psi}(\mathbb{R}^{2})$.
Proof.
From Lemma 2.1 and [17, Theorem I.III.3], there holds
$\Psi\bigl(\Phi^{\prime}(t)\bigr)=t\Phi^{\prime}(t)-\Phi(t)\leq(\kappa-1)\Phi(t).\qed$
We conclude with the following properties about convergence and boundedness.
Lemma 2.3.
Consider $(u_{n})\subset L^{\Phi}$.
  • Let $u\in L^{\Phi}$. If $\lim_{n}|u_{n}-u|_{\Phi}=0$, then $\lim_{n}\int_{\mathbb{R}^{2}}\Phi(u_{n}-u)\,dx=0$. The opposite implication holds if $\Phi$ satisfies the $\Delta_{2}$ condition globally.
  • Assume that $\Phi$ satisfies the $\Delta_{2}$ condition globally. Then, $(u_{n})$ is bounded if and only if $\left(\int_{\mathbb{R}^{2}}\Phi(|u_{n}|)\,dx\right)$ is bounded.
Let us introduce the space
${\mathcal{V}}$ $:=$ $\Big\{u=\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in H^{1}(\mathbb{R}^{2})^{6}:\Big\langle u,\overset{\circ}{\nabla}\begin{pmatrix}\alpha\\ \widetilde{\alpha}\end{pmatrix}\Big\rangle_{2}=0\mbox{ for any }\alpha,\widetilde{\alpha}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})\Big\}$
$=$ $\Big\{u=\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in H^{1}(\mathbb{R}^{2})^{6}:\partial_{x_{1}}u_{1}+\partial_{x_{2}}u_{2}+k\widetilde{u}_{3}=0,\partial_{x_{1}}\widetilde{u}_{1}+\partial_{x_{2}}\widetilde{u}_{2}-ku_{3}=0\mbox{ a.e. in }\mathbb{R}^{2}\Big\},$
and note that it is a closed subspace of $H^{1}(\mathbb{R}^{2})^{6}$. Let us consider the norm in $\mathcal{V}$}
$\|u\|:=\Big(\sum_{i=1}^{3}|\nabla u_{i}|^{2}_{2}+k^{2}|u_{i}|^{2}_{2}+|\nabla\widetilde{u}_{i}|^{2}_{2}+k^{2}|\widetilde{u}_{i}|^{2}_{2}\Big)^{1/2},$
which is equivalent to the standard $H^{1}$-norm. Let $\mathcal{W}$} be the completion of vector fields $w=\overset{\circ}{\nabla}\begin{pmatrix}\phi\\ \widetilde{\phi}\end{pmatrix}$, where $\phi,\widetilde{\phi}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{3}$, with respect to the norm
$\|w\|:=\big(|w|_{2}^{2}+|w|_{\Phi}^{2}\big)^{1/2}$
so that ${\mathcal{W}}\subset L^{2}(\mathbb{R}^{2})^{6}\cap L^{\Phi}$. We will see that ${\mathcal{V}}\cap{\mathcal{W}}=\{0\}$ (cf. Theorem 2.4 below), and we define a norm on the space
$X:={\mathcal{V}}\oplus{\mathcal{W}}$
as
$\|v+w\|^{2}=\|v\|^{2}+\|w\|^{2},\quad v\in{\mathcal{V}},w\in{\mathcal{W}}.$
If $\Phi$ is the N-function in (F2)–(F3), then, from the $\Delta_{2}$ condition – cf. (N1) – and Lemma 2.1, it has at most a power-like growth at infinity. Taking (N3) into account, there exists $p>2$ such that
$\lim_{|t|\to\infty}\frac{|\Phi^{\prime}(t)|}{|t|^{p-1}}=0.$
Then, using also (N2), for every $t\in\mathbb{R}$ we have $\Phi(t)\lesssim|t|^{2}+|t|^{p}$, thus
${\mathcal{V}}\subset H^{1}(\mathbb{R}^{2})^{6}\subset L^{2}(\mathbb{R}^{2})^{6}\cap L^{p}(\mathbb{R}^{2})^{6}\subset L^{\Phi}.$
Moreover, using Lemma 2.2 and arguing as in [8, Lemma 2.1], we can prove that $J\in{\mathcal{C}}^{1}(X)$.
Next, again from the $\Delta_{2}$ condition and Lemma 2.1, $\Phi$ has at most a power-type behaviour at the origin. Taking (N2) into account, there exists $q>2$ such that
$\lim_{t\to 0}\frac{|\Phi^{\prime}(t)|}{|t|^{q-1}}=\infty.$
Then, using also (N3) and since we can assume $q=p$, we have $|t|^{p}\lesssim\Phi(t)$ for every $t\in[-1,1]$ and $|t|^{2}\lesssim\Phi(t)$ for every $t\in(-\infty,-1)\cup(1,\infty)$, thus
$L^{\Phi}\subset L^{2}(\mathbb{R}^{2})^{6}+L^{p}(\mathbb{R}^{2})^{6}.$
Exploiting Lemma 2.1 once again, we infer
$L^{2}(\mathbb{R}^{2})^{6}\cap L^{p^{\prime}}(\mathbb{R}^{2})^{6}\subset L^{\Psi},$
where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. This will be used in the proof of Lemma 2.4 below.
Recall that the second-order differential operator $L\colon\operatorname{dom}(L)\subset L^{2}(\mathbb{R}^{2})^{6}\to L^{2}(\mathbb{R}^{3})^{6}$ is elliptic and self-adjoint on the domain $\operatorname{dom}(L)=\{u\in L^{2}(\mathbb{R}^{2})^{6}:Lu\in L^{2}(\mathbb{R}^{2})^{6}\}$. Its associated bilinear form $b_{L}:\operatorname{dom}(b_{L})\times\operatorname{dom}(b_{L})\to\mathbb{R}$ is given by
$b_{L}(u,v)=\int_{\mathbb{R}^{2}}\overset{\circ}{\nabla}\times u\cdot\overset{\circ}{\nabla}\times v\,dx$
with $\operatorname{dom}(b_{L})=\{u\in L^{2}(\mathbb{R}^{2})^{6}:\overset{\circ}{\nabla}\times u\in L^{2}(\mathbb{R}^{2})^{6}\}$. Recall also that the operator $L\colon\operatorname{dom}(L)\subset L^{2}(\mathbb{R}^{2})^{6}\to L^{2}(\mathbb{R}^{3})^{6}$ has spectrum $\sigma(L)=\{0\}\cup[k^{2},\infty)$, where 0 is an eigenvalue of infinite multiplicity and $k^{2},\infty)$ consists of absolutely continuous spectrum. Finally,
$b_{L}(v,v)=\|v\|^{2},\quad\hbox{for }v\in{\mathcal{V}}.$
Theorem 2.4.
The spaces $\mathcal{V}$} and $\mathcal{W}$} are closed subspaces of $L^{2}(\mathbb{R}^{2})^{6}$ and orthogonal with respect to $\langle\cdot,\cdot\rangle_{2}$, and $X={\mathcal{V}}\oplus{\mathcal{W}}$ is the completion of ${\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ with respect to the norm $\|\cdot\|$. Moreover, $\mathrm{cl\,}_{L^{\Phi}}{\mathcal{V}}\cap\mathrm{cl\,}_{L^{\Phi}}{\mathcal{W}}=\{0\}$.
Proof.
In order to get the Helmholtz decomposition we argue in a similar way as in [14, Theorem 2.1], and for the reader’s convenience we sketch this part, which will be useful to prove also the latter statement that $\mathrm{cl\,}_{L^{\Phi}}{\mathcal{V}}\cap\mathrm{cl\,}_{L^{\Phi}}{\mathcal{W}}=\{0\}$.
Note that $\mathcal{V}$} and $\mathcal{W}$} are closed subspaces of $L^{2}(\mathbb{R}^{2})^{6}$ and orthogonal with respect to $\langle\cdot,\cdot\rangle_{2}$. Let $\varphi=\begin{pmatrix}\phi\\ \widetilde{\phi}\end{pmatrix}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ and let $\alpha,\widetilde{\alpha}\in H^{1}(\mathbb{R}^{2})\cap{\mathcal{C}}^{\infty}(\mathbb{R}^{2})$ be the unique solutions to
$\Delta\alpha+k^{2}\alpha=-\big(\partial_{x_{1}}\phi_{1}+\partial_{x_{2}}\phi_{2}+k\widetilde{\phi}_{3}\big)$
and
$\Delta\widetilde{\alpha}+k^{2}\widetilde{\alpha}=-\big(\partial_{x_{1}}\widetilde{\phi}_{1}+\partial_{x_{2}}\widetilde{\phi}_{2}-k\phi_{3}\big)$
respectively. Since $\alpha$ is the Bessel potential of the ${\mathcal{C}}_{0}^{\infty}$-function $(\partial_{x_{1}}\phi_{1}+\partial_{x_{2}}\phi_{2}+k\widetilde{\phi}_{3})$ we find that $\alpha\in W^{l,s}(\mathbb{R}^{2})$ for every $l\in\mathbb{N}$ and $1<s<\infty$, cf. [28], and similarly for $\widetilde{\alpha}$. Moreover, we find $\alpha_{n},\widetilde{\alpha}_{n}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})$ such that $\alpha_{n}\to\alpha$ and $\widetilde{\alpha}_{n}\to\widetilde{\alpha}$ in $W^{l,s}(\mathbb{R}^{2})$ for any $l\in\mathbb{N}$ and $1<s<\infty$. In particular, this implies
$\varphi_{\mathcal{W}}:=\overset{\circ}{\nabla}\begin{pmatrix}\alpha\\ \widetilde{\alpha}\end{pmatrix}=\lim_{n\to\infty}\overset{\circ}{\nabla}\begin{pmatrix}\alpha_{n}\\ \widetilde{\alpha}_{n}\end{pmatrix}\in{\mathcal{W}},$
and $\varphi_{\mathcal{W}}\in W^{l,s}(\mathbb{R}^{2})$ for all $l\in\mathbb{N}$ and $1<s<\infty$. In addition, $\mathcal{W}$} is clearly contained in the completion of ${\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ with respect to the norm $\|\cdot\|$. Since, in particular, $\varphi_{\mathcal{W}}\in H^{1}(\mathbb{R}^{2})^{6}$, we obtain by integration by parts
$\Big\langle\varphi,\overset{\circ}{\nabla}\begin{pmatrix}\beta\\ \widetilde{\beta}\end{pmatrix}\Big\rangle_{2}$ $=$ $-\int_{\mathbb{R}^{2}}(\partial_{x_{1}}\phi_{1}+\partial_{x_{2}}\phi_{2}+k\widetilde{\phi}_{3})\beta+(\partial_{x_{1}}\widetilde{\phi}_{1}+\partial_{x_{2}}\widetilde{\phi}_{2}-k\phi_{3})\widetilde{\beta}\,dx$
$=$ $\int_{\mathbb{R}^{2}}(-\Delta\alpha+k^{2}\alpha)\beta+(-\Delta\widetilde{\alpha}+k^{2}\widetilde{\alpha})\widetilde{\beta}\,dx=\Big\langle\varphi_{\mathcal{W}},\overset{\circ}{\nabla}\begin{pmatrix}\beta\\ \widetilde{\beta}\end{pmatrix}\Big\rangle_{2}$
for every $\beta,\widetilde{\beta}\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})$. Therefore, $\varphi_{\mathcal{V}}:=\varphi-\varphi_{\mathcal{W}}\in{\mathcal{V}}$, and we obtain the following Helmholtz-type decomposition
[…]
Moreover, we have shown that $\varphi\in X$. It remains to show that also $\mathcal{V}$} is contained in the completion of ${\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ with respect to $\|\cdot\|$. To see this, let $v\in{\mathcal{V}}$ and take $(\varphi_{n})\subset{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ such that $\varphi_{n}\to v$ in $H^{1}(\mathbb{R}^{2})^{6}$. Let us decompose $\varphi_{n}=\varphi_{\mathcal{V}}^{n}+\varphi_{\mathcal{W}}^{n}\in{\mathcal{V}}\oplus{\mathcal{W}}$ and let $\mathrm{cl\,}_{L^{2}\cap L^{\Phi}}{\mathcal{V}}$ denote the closure of $\mathcal{V}$} in $L^{2}(\mathbb{R}^{2})^{6}\cap L^{\Phi}$. Observe that $\big(\mathrm{cl\,}_{L^{2}\cap L^{\Phi}}{\mathcal{V}}\big)\cap{\mathcal{W}}=\{0\}$, so there is a continuous $L^{2}(\mathbb{R}^{2})^{6}\cap L^{\Phi}$-projection of $\big(\mathrm{cl\,}_{L^{2}\cap L^{\Phi}}{\mathcal{V}}\big)\oplus{\mathcal{W}}$ onto $\mathcal{W}$}. Since $\varphi_{n}\to v$ in $L^{2}(\mathbb{R}^{2})^{6}\cap L^{\Phi}$, we infer that $\varphi_{\mathcal{W}}^{n}\to 0$ in $\mathcal{W}$}. Similarly, arguing with the closure of $\mathcal{W}$} in $H^{1}(\mathbb{R}^{2})^{6}$ we get $\varphi_{\mathcal{V}}^{n}\to v$ in $\mathcal{V}$}. Therefore
$\|v-\varphi_{n}\|^{2}=\|v-\varphi_{\mathcal{V}}^{n}\|^{2}+\|\varphi^{n}_{\mathcal{W}}\|^{2}\to 0$
as $n\to\infty$, and we conclude that ${\mathcal{V}}\oplus{\mathcal{W}}$ is the completion of ${\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$ with respect to the norm $\|\cdot\|$.
Now, let $u\in\mathrm{cl\,}_{L^{\Phi}}{\mathcal{V}}\cap\mathrm{cl\,}_{L^{\Phi}}{\mathcal{W}}$ and $\varphi=\varphi_{\mathcal{V}}+\varphi_{\mathcal{W}}$ as above. Consider $(\varphi_{n})\subset{\mathcal{V}}$ and $(\psi_{n})\subset{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})$ such that $\lim_{n}|u-\varphi_{n}|_{\Phi}=\lim_{n}|u-\overset{\circ}{\nabla}\psi_{n}|_{\Phi}=0$ as $n\to\infty$. Observe that $\varphi_{\mathcal{V}},\varphi_{\mathcal{W}}\in L^{2}(\mathbb{R}^{2})^{6}\cap L^{p^{\prime}}(\mathbb{R}^{2})^{6}\subset L^{\Psi}$, whence
$\int_{\mathbb{R}^{2}}u\varphi_{\mathcal{W}}\,dx$ $=\lim_{n\to\infty}\int_{\mathbb{R}^{2}}\varphi_{n}\overset{\circ}{\nabla}\varphi\,dx=0,$
$\int_{\mathbb{R}^{2}}u\varphi_{\mathcal{V}}\,dx$ $=\lim_{n\to\infty}\int_{\mathbb{R}^{2}}\overset{\circ}{\nabla}\psi_{n}\varphi_{\mathcal{V}}\,dx=0.$
Therefore, $\int_{\mathbb{R}^{2}}u\varphi\,dx=0$ for every $\varphi\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$, and $u=0$. ∎
3. Analysis of Cerami sequences
Note that (F1) and (F2) imply that for any $\varepsilon>0$ there is $C_{\varepsilon}>0$ such that
(3.1) $F(u)\leq\varepsilon|u|^{2}+C_{\varepsilon}\Phi(u)\quad\text{for every }u\in\mathbb{R}^{6}.$
For $v\in{\mathcal{V}}$, let $w(v)$ denote the unique minimizer of
$w\mapsto\frac{1}{2}\int_{\mathbb{R}^{2}}V(x)|v+w|^{2}\,dx+\int_{\mathbb{R}^{2}}F(v+w)\,dx$
in ${\mathcal{W}}\subset L^{2}(\mathbb{R}^{2})^{6}\cap L^{\Phi}$. Let $\widetilde{J}\colon{\mathcal{V}}\to\mathbb{R}$ be given by
$\widetilde{J}(v):=J(v+w(v))=\frac{1}{2}\big(|\nabla v|^{2}_{2}+k^{2}|v|^{2}_{2}\big)-\Big(\frac{1}{2}\int_{\mathbb{R}^{2}}V(x)|v+w(v)|^{2}\,dx+\int_{\mathbb{R}^{2}}F(v+w(v))\,dx\Big).$
In a way similar to [6, Proof of Theorem 4.4] we show that $\widetilde{J}\in{\mathcal{C}}^{1}({\mathcal{V}})$ and
$\widetilde{J}^{\prime}(v)(\varphi)=J^{\prime}(v+w(v))(\varphi)$
for every $v,\varphi\in{\mathcal{V}}$. In view of (V) and (3.1), we find $r>0$ such that
(3.2) $\inf_{\|v\|=r,\,v\in{\mathcal{V}}}\widetilde{J}(v)\geq\inf_{\|v\|=r,\,v\in{\mathcal{V}}}J(v)>0.$
Arguing as in the proof of [14, Lemma 6.3] and using Lemma 2.2 we obtain the following property.
Lemma 3.1.
If (F3) is satisfied, then for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that for every $u=v+w\in{\mathcal{V}}\oplus{\mathcal{W}}$ and every $t\in[0,\sqrt{1-2/\gamma}]$
$J(u)\geq J(tv)+J^{\prime}(u)\biggl(\frac{1-t^{2}}{2}u+t^{2}w\biggr)-\varepsilon t^{2}|u|_{2}|w|_{2}-t^{2}c_{\varepsilon}\big|\Phi^{\prime}(|u|)\big|_{\Psi}|w|_{\Phi}.$
Lemma 3.2.
If $v\in{\mathcal{V}}$ is such that $\overset{\circ}{\nabla}\times v=0$ in the sense of distributions, then $v=0$.
Proof.
In this proof, we shall denote points in $\mathbb{R}^{3}$ by $(x,y,z)$. Let $v=\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}$ and define $E$ via (1.1) with $t=0$. Observe that $\overset{\circ}{\nabla}\times v=0$ implies $\nabla\times E=0$, and $v\in{\mathcal{V}}$ implies $\mathop{\mathrm{div}\,}E=0$, both in the sense of distributions. From $\nabla\times E=0$, [10, Lemma 1.1 (i)], and the fact that $E\in L^{2}_{\textup{loc}}(\mathbb{R}^{3})$, there exists $\xi\in H^{1}_{\textup{loc}}(\mathbb{R}^{3})$ such that $E=\nabla\xi$. Then, $0=\mathop{\mathrm{div}\,}E=\Delta\xi$, so $E$ is harmonic in $\mathbb{R}^{3}$ as well. Writing $\Delta E(x,y,z)=0$ explicitly, we obtain
$\begin{split}\partial_{x}^{2}v(x,y)\cos(kz)&+\partial_{x}^{2}\widetilde{v}(x,y)\sin(kz)+\partial_{y}^{2}v(x,y)\cos(kz)+\partial_{y}^{2}\widetilde{v}(x,y)\sin(kz)\\ &-k^{2}v(x,y)\cos(kz)-k^{2}\widetilde{v}(x,y)\sin(kz)=0,\end{split}$
whence, taking $z=0$ and $z=\pi/(2k)$,
$\Delta v+k^{2}v=-\Delta\widetilde{v}+k^{2}\widetilde{v}=0\quad\text{in }\mathbb{R}^{2}.$
Since $v\in H^{1}(\mathbb{R}^{2})$, we conclude. ∎
The next Lemma refines the abstract condition [15, (I7)].
Lemma 3.3.
Let $(v_{n})\subset{\mathcal{V}}$, $v\in{\mathcal{V}}\setminus\{0\}$, $(w_{n})\subset{\mathcal{W}}$, and $(t_{n})\subset(0,\infty)$ such that $t_{n}\to\infty$ and $v_{n}\rightharpoonup v$ as $n\to\infty$. Then
$\lim_{n}\frac{1}{t_{n}^{2}}\int_{\mathbb{R}^{2}}F\bigl(t_{n}(v_{n}+w_{n})\bigr)\,dx=\infty.$
Proof.
Observe that (F3) implies
$\frac{1}{t_{n}^{2}}\int_{\mathbb{R}^{2}}F\bigl(t_{n}(v_{n}+w_{n})\bigr)\,dx\geq\frac{c_{2}}{t_{n}^{2}}\int_{\mathbb{R}^{2}}\Phi\bigl(t_{n}|v_{n}+w_{n}|\bigr)\,dx.$
From (N3), there exist $C,D>0$ such that for every $t\in\mathbb{R}$, $C\Phi(t)\geq t^{2}-D$. For all $R>0$,
$\int_{B(0,R)}|v_{n}+w_{n}|^{2}\,dx\leq\frac{C}{t_{n}^{2}}\int_{\mathbb{R}^{2}}\Phi\bigl(t_{n}|v_{n}+w_{n}|\bigr)\,dx+\pi R^{2}D,$
thus the statement is true if $|v_{n}+w_{n}|$ is unbounded in $L^{2}(B(0,R))$ for some $R>0$.
Now, assume that $|v_{n}+w_{n}|$ is bounded in $L^{2}(B(0,R))$ for all $R>0$. Up to a subsequence, $v_{n}(x)\to v(x)$ for a.e. $x\in\mathbb{R}^{2}$ and $w_{n}\to w$ in $L^{2}_{\textup{loc}}(\mathbb{R}^{2})$ for some $w$ as $n\to\infty$. For every $\varepsilon>0$ and every $n$, we define
$\Omega_{n}^{\varepsilon}:=\left\{x\in\mathbb{R}^{2}:|v_{n}(x)+w_{n}(x)|\geq\varepsilon\right\}.$
Let us assume by contradiction that $\lim_{n}|\Omega_{n}^{\varepsilon}|=0$ for all $\varepsilon>0$, i.e. $\lim_{n}|v_{n}+w_{n}|=0$ in measure. Then, up to a subsequence, $v_{n}(x)+w_{n}(x)\to 0$ for a.e. $x\in\mathbb{R}^{2}$ as $n\to\infty$. Consequently, $w_{n}\to-v$ a.e. in $\mathbb{R}^{2}$ and, therefore, $w_{n}\rightharpoonup-v$ in $L^{2}_{\textup{loc}}(\mathbb{R}^{2})$ as $n\to\infty$. Since $\overset{\circ}{\nabla}\times w_{n}=0$ for every $n$ in the sense of distributions, this implies that the same holds for $v$. From Lemma 3.2, $v=0$, a contradiction. Finally, taking $\varepsilon>0$ such that $\lim_{n}|\Omega_{n}^{\varepsilon}|>0$ along a subsequence, we get
$\lim_{n}\int_{\mathbb{R}^{2}}\frac{\Phi\bigl(t_{n}|v_{n}+w_{n}|\bigr)}{t_{n}^{2}}\,dx\geq\lim_{n}\int_{\Omega_{n}^{\varepsilon}}\frac{\Phi\bigl(t_{n}|v_{n}+w_{n}|\bigr)}{t_{n}^{2}|v_{n}+w_{n}|}|v_{n}+w_{n}|\,dx=\infty.\qed$
Lemma 3.4.
If $X\subset{\mathcal{V}}$ has a finite dimension and $(v_{n})\subset X$ is such that $\lim_{n}\|v_{n}\|=\infty$, then $\lim_{n}\widetilde{J}(v_{n})=$-$\infty$.
Proof.
Let us denote $\widetilde{v}_{n}:=v_{n}/\|v_{n}\|$. Up to a subsequence, there exists $\widetilde{v}\in X$ such that $\lim_{n}\widetilde{v}=v$. In particular, $\widetilde{v}\neq 0$. Using Lemma 3.3, we obtain
$\lim_{n}\frac{\widetilde{J}(v_{n})}{\|v_{n}\|^{2}}\leq\frac{1}{2}-\frac{1}{\|v_{n}\|^{2}}\int_{\mathbb{R}^{2}}F\bigl(\|v_{n}\|\widetilde{v}_{n}+w(v_{n})\bigr)\,dx=-\infty.\qed$
Normally, the Ambrosetti–Rabinowitz-type condition (F3) (cf. [2]) provides the boundedness of Palais–Smale sequences by simply estimating $\widetilde{J}(v_{n})-\frac{1}{\gamma}\widetilde{J}^{\prime}(v_{n})(v_{n})$ . As we shall see below, in our problem obtaining the boundedness is much more complicated even if $F(u)=\frac{1}{p}|u|^{p}$, and achieved only for Cerami sequences (cf. [7]).
Lemma 3.5.
If $(v_{n})\subset{\mathcal{V}}$ is a Cerami sequence for $\widetilde{J}$ such that $\liminf_{n}\widetilde{J}(v_{n})\geq 0$, then it is bounded.
We recall that a Cerami sequence for $\widetilde{J}$ is a sequence $(v_{n})\subset{\mathcal{V}}$ such that $\bigl(\widetilde{J}(v_{n})\bigr)$ is bounded and $\lim_{n}(1+\|v_{n}\|)\widetilde{J}^{\prime}(v_{n})=0$, and that a Cerami sequence at the level $c\in\mathbb{R}$ is a Cerami sequence such that $\lim_{n}\widetilde{J}(v_{n})=c$.
Proof of Lemma 3.5.
Let $\beta\geq\limsup_{n}\widetilde{J}(v_{n})$. For every $n$, let us denote $w_{n}:=w(v_{n})$ and $u_{n}:=v_{n}+w_{n}$. From the second inequality in (F3),
$\frac{1}{2}\|v_{n}\|^{2}-\frac{1}{2}\operatorname*{ess\,inf}V|v_{n}|_{2}^{2}\geq J(u_{n})+\frac{1}{2}\operatorname*{ess\,inf}V|w_{n}|_{2}^{2}+c_{2}\int_{\mathbb{R}^{2}}\Phi(|u_{n}|)\,dx.$
Adding $\|v_{n}\|^{2}/2-\operatorname*{ess\,inf}V|v_{n}|_{2}^{2}/2$ on both sides, we obtain
(3.3) $\|v_{n}\|^{2}-\operatorname*{ess\,inf}V|v_{n}|_{2}^{2}\\ \geq J(u_{n})+\frac{1}{2}\|v_{n}\|^{2}-\frac{1}{2}\operatorname*{ess\,inf}V|v_{n}|_{2}^{2}+\frac{1}{2}\operatorname*{ess\,inf}V|w_{n}|_{2}^{2}+c_{2}\int_{\mathbb{R}^{2}}\Phi(|u_{n}|)\,dx.$
Set
$s_{n}^{2}:=\frac{1}{2}\|v_{n}\|^{2}-\frac{1}{2}\operatorname*{ess\,inf}V|v_{n}|_{2}^{2}+\frac{1}{2}\operatorname*{ess\,inf}V|w_{n}|_{2}^{2}+c_{2}\int_{\mathbb{R}^{2}}\Phi(|u_{n}|)\,dx$
and $\widetilde{v}_{n}:=v_{n}/s_{n}$. Assume by contradiction that $(u_{n})$ is unbounded. From Theorem 2.4, this implies that $\lim_{n}s_{n}=\infty$ along a subsequence. We want to prove that
(3.4) $\liminf_{n}|\widetilde{v}_{n}|_{\Phi}>0,$
thus let us assume by contradiction that $|\widetilde{v}_{n}|_{\Phi}\to 0$ along a subsequence as $n\to\infty$. This, together with (F1) and (F2), implies that for every $s>0$
(3.5) $\lim_{n}\int_{\mathbb{R}^{2}}F(s\widetilde{v}_{n})\,dx=0.$
From (F3), there holds
$J(u_{n})-\frac{1}{2}J^{\prime}(u_{n})(u_{n})\geq\frac{\gamma-2}{2}\int_{\mathbb{R}^{2}}F(u_{n})\,dx,$
thus $(u_{n})$ is bounded in $L^{\Phi}$. From Theorem 2.4, $(w_{n})$ is bounded in $L^{\Phi}$. Moreover, there exists $\delta>0$ such that $s_{n}\geq\delta|u_{n}|_{2}^{2}\geq\delta|w_{n}|_{2}^{2}$. From Lemma 3.1 with $u=u_{n}$ and $t=s/s_{n}$ (observe that $t\leq\sqrt{1-2/\gamma}$ for $n\gg 1$) and Lemma 2.2, for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\begin{split}\beta&\geq\liminf_{n}J(u_{n})\\ &\geq\liminf_{n}J(s\widetilde{v}_{n})+\liminf_{n}J^{\prime}(u_{n})\biggl(\frac{1-(s/s_{n})^{2}}{2}u_{n}+(s/s_{n})^{2}w_{n}\biggr)\\ &\quad-\varepsilon\limsup_{n}(s/s_{n})^{2}|u_{n}|_{2}|w_{n}|_{2}-c_{\varepsilon}\limsup_{n}(s/s_{n})^{2}\big|\Phi^{\prime}(|u_{n}|)\big|_{\Psi}|w_{n}|_{\Phi}\\ &\geq\liminf_{n}J(s\widetilde{v}_{n})+\liminf_{n}J^{\prime}(u_{n})\biggl(\frac{1-(s/s_{n})^{2}}{2}u_{n}+(s/s_{n})^{2}w_{n}\biggr)-s^{2}\frac{\varepsilon}{\delta}.\end{split}$
Observe that the minimality of $w_{n}$ implies $J^{\prime}(u_{n})(w_{n})=0$, hence, since $\lim_{n}J^{\prime}(u_{n})(u_{n})=0$, we obtain
$\beta\geq\liminf_{n}J(u_{n})-s^{2}\frac{\varepsilon}{\delta},$
which, together with (3.5), yields
$\beta\geq s^{2}\biggl(\liminf_{n}\bigl(\|\widetilde{v}_{n}\|^{2}-\operatorname*{ess\,sup}V|\widetilde{v}_{n}|_{2}^{2}\bigr)-\frac{\varepsilon}{\delta^{2}}\biggr).$
Taking $\sigma\in(0,1)$ such that
$\operatorname*{ess\,sup}V\leq\frac{\sigma}{2}\operatorname*{ess\,inf}V+(1-\sigma)k^{2},$
we obtain, for every $n$,
$\begin{split}\|\widetilde{v}_{n}\|^{2}-\operatorname*{ess\,sup}V|\widetilde{v}_{n}|_{2}^{2}&\geq\|\widetilde{v}_{n}\|^{2}-(1-\sigma)k^{2}|\widetilde{v}_{n}|_{2}^{2}-\frac{\sigma}{2}\operatorname*{ess\,inf}V|\widetilde{v}_{n}|_{2}^{2}\\ &\geq\sigma\biggl(\|\widetilde{v}_{n}\|^{2}-\frac{1}{2}\operatorname*{ess\,inf}V|\widetilde{v}_{n}|_{2}^{2}\biggr)\geq\sigma+o_{n}(1),\end{split}$
where the last inequality comes from (3.3) and $J(u_{n})\geq o_{n}(1)$. In conclusion, taking $s=2\sqrt{\beta/\sigma}$ and $\varepsilon=\sigma\delta^{2}/8$, we get the contradiction
$\beta\geq s^{2}\biggl(\frac{\sigma}{2}-\frac{\varepsilon}{\delta^{2}}\biggr)=\frac{3}{2}\beta.$
Using (3.4) and the fact that $(\widetilde{v}_{n})$ is bounded, we find $\widetilde{v}\in{\mathcal{V}}\setminus\{0\}$ such that $\widetilde{v}_{n}\rightharpoonup\widetilde{v}$ as $n\to\infty$ along a subsequence. Then, from Lemma 3.3,
$0=\lim_{n}\frac{J(u_{n})}{s_{n}^{2}}\leq\lim_{n}\frac{1}{2}\|\widetilde{v}_{n}\|^{2}-\frac{1}{s_{n}^{2}}\int_{\mathbb{R}^{2}}F(s_{n}\widetilde{v}_{n}+w_{n})\,dx=-\infty,$
a contradiction. ∎
Now we introduce the cylindrical symmetry of the problem. Let
$g=\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{pmatrix}\in{\mathcal{S}}{\mathcal{O}}(2)$
with $\alpha\in\mathbb{R}$, and let $U\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$. We define
$(g\star U)(x):=\begin{pmatrix}g&0\\ 0&1\end{pmatrix}U(g^{-1}x)=\begin{pmatrix}U_{1}(g^{-1}x)\cos\alpha-U_{2}(g^{-1}x)\sin\alpha\\ U_{1}(g^{-1}x)\sin\alpha+U_{2}(g^{-1}x)\cos\alpha\\ U_{3}(g^{-1}x)\end{pmatrix}$
for $x\in\mathbb{R}^{2}$, where $g^{-1}x=(x_{1}\cos\alpha+x_{2}\sin\alpha,-x_{1}\sin\alpha+x_{2}\cos\alpha)$.
Lemma 3.6.
There holds
[…]
for $U,\widetilde{U}\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ and $g\in{\mathcal{S}}{\mathcal{O}}(2)$.
Proof.
In the following, when we write $\partial_{x_{i}}U_{j}(g^{-1}x)$ we always mean $(\partial_{x_{i}}U_{j})(g^{-1}x)$.
Observe that
$\overset{\circ}{\nabla}\times\begin{pmatrix}g\star U\\ g\star\widetilde{U}\end{pmatrix}$ $:=$ $\begin{pmatrix}\partial_{x_{1}}U_{3}(g^{-1}x)\sin\alpha+\partial_{x_{2}}U_{3}(g^{-1}x)\cos\alpha-k\widetilde{U}_{1}(g^{-1}x)\sin\alpha-k\widetilde{U}_{2}(g^{-1}x)\cos\alpha\\ k\widetilde{U}_{1}(g^{-1}x)\cos\alpha-k\widetilde{U}_{2}(g^{-1}x)\sin\alpha-\partial_{x_{1}}U_{3}(g^{-1}x)\cos\alpha+\partial_{x_{2}}U_{3}(g^{-1}x)\sin\alpha\\ \partial_{x_{1}}U_{2}(g^{-1}x)-\partial_{x_{2}}U_{1}(g^{-1}x)\\ \partial_{x_{1}}\widetilde{U}_{3}(g^{-1}x)\sin\alpha+\partial_{x_{2}}\widetilde{U}_{3}(g^{-1}x)\cos\alpha+kU_{1}(g^{-1}x)\sin\alpha+kU_{2}(g^{-1}x)\cos\alpha\\ -kU_{1}(g^{-1}x)\cos\alpha+kU_{2}(g^{-1}x)\sin\alpha-\partial_{x_{1}}\widetilde{U}_{3}(g^{-1}x)\cos\alpha+\partial_{x_{2}}\widetilde{U}_{3}(g^{-1}x)\sin\alpha\\ \partial_{x_{1}}\widetilde{U}_{2}(g^{-1}x)-\partial_{x_{2}}\widetilde{U}_{1}(g^{-1}x)\end{pmatrix}$
$=$ $\begin{pmatrix}\partial_{x_{1}}U_{3}(g^{-1}x)\sin\alpha+\partial_{x_{2}}U_{3}(g^{-1}x)\cos\alpha-k\widetilde{U}_{1}(g^{-1}x)\sin\alpha-k\widetilde{U}_{2}(g^{-1}x)\cos\alpha\\ -\partial_{x_{1}}U_{3}(g^{-1}x)\cos\alpha+\partial_{x_{2}}U_{3}(g^{-1}x)\sin\alpha+k\widetilde{U}_{1}(g^{-1}x)\cos\alpha-k\widetilde{U}_{2}(g^{-1}x)\sin\alpha\\ \partial_{x_{1}}U_{2}(g^{-1}x)-\partial_{x_{2}}U_{1}(g^{-1}x)\\ \partial_{x_{1}}\widetilde{U}_{3}(g^{-1}x)\sin\alpha+\partial_{x_{2}}\widetilde{U}_{3}(g^{-1}x)\cos\alpha+kU_{1}(g^{-1}x)\sin\alpha+kU_{2}(g^{-1}x)\cos\alpha\\ -\partial_{x_{1}}\widetilde{U}_{3}(g^{-1}x)\cos\alpha+\partial_{x_{2}}\widetilde{U}_{3}(g^{-1}x)\sin\alpha-kU_{1}(g^{-1}x)\cos\alpha+kU_{2}(g^{-1}x)\sin\alpha\\ \partial_{x_{1}}\widetilde{U}_{2}(g^{-1}x)-\partial_{x_{2}}\widetilde{U}_{1}(g^{-1}x)\end{pmatrix}$
and we easily conclude that
$\Big|\overset{\circ}{\nabla}\times\begin{pmatrix}(g\star U)(x)\\ (g\star\widetilde{U})(x)\end{pmatrix}\Big|^{2}=\Big|\overset{\circ}{\nabla}\times\begin{pmatrix}U(g^{-1}x)\\ \widetilde{U}(g^{-1}x)\end{pmatrix}\Big|^{2}.\qed$
Let
${\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ $:=$ $\big\{v\in{\mathcal{V}}:\;g\star v=v\hbox{ for any }g\in{\mathcal{S}}{\mathcal{O}}(2)\big\},$
${\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ $:=$ $\big\{w\in{\mathcal{W}}:\;g\star w=w\hbox{ for any }g\in{\mathcal{S}}{\mathcal{O}}(2)\big\}.$
From now on, for $v\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, $w(v)$ is the unique minimizer of
$w\mapsto\frac{1}{2}\int_{\mathbb{R}^{2}}V(x)|v+w|^{2}\,dx+\int_{\mathbb{R}^{2}}F(v+w)\,dx$
in ${\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$.
Lemma 3.7.
The functional $\widetilde{J}|_{{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}}$ satisfies the Cerami condition at every nonngative level (that is, every Cerami sequence at a nonnegative level has a converging subsequence).
Proof.
First, we claim that ${\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ is compactly embedded into $L^{r}(\mathbb{R}^{2})^{6}$ for any $r>2$. Indeed, let $(v_{n})\subset{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ be such that $v_{n}\rightharpoonup 0$ as $n\to\infty$. Passing to a subsequence $v_{n}\to 0$ a.e. on $\mathbb{R}^{2}$ as $n\to\infty$. Then, $(|v_{n}|)$ is bounded in $H^{1}(\mathbb{R}^{2})$ and ${\mathcal{O}}(2)$-invariant. Hence, $|v_{n}|\to v_{0}$ in $L^{r}(\mathbb{R}^{2})$ along a subsequence as $n\to\infty$, which yields $v_{0}=0$.
Next, let $(v_{n})\subset{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ be a Cerami sequence for $\widetilde{J}|_{{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}}$ as in the statement. From Lemma 3.5, there exists $v\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ such that $v_{n}\rightharpoonup v$ along a subsequence as $n\to\infty$. In particular, up to taking other subsequences, $v_{n}\to v$ in $L^{\Phi}$ and a.e. in $\mathbb{R}^{2}$ as $n\to\infty$. From the continuity of $w\colon L^{\Phi}_{{\mathcal{S}}{\mathcal{O}}(2)}\to{\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, which is proved as in [15, Lemma 4.6], $(w_{n})\subset{\mathcal{W}}$ is bounded, where $w_{n}:=w(v_{n})$. There holds
$\begin{split}o_{n}(1)&=\widetilde{J}^{\prime}(v_{n})(v_{n})-\widetilde{J}^{\prime}(v_{n})(v)\\ &=b_{L}(v_{n},v_{n}-v)-\int_{\mathbb{R}^{2}}V(x)\langle v_{n}+w_{n},v_{n}-v\rangle+\langle f(v_{n}+w_{n}),v_{n}-v\rangle\,dx\\ &=b_{L}(v_{n},v_{n}-v)-\int_{\mathbb{R}^{2}}V(x)\langle v_{n},v_{n}-v\rangle\,dx+o_{n}(1),\end{split}$
which implies that $v_{n}\to v$ as $n\to\infty$. ∎
Remark 3.8.
Our proof can be also adopted in the situation where the Ambrosetti–Rabinowitz-type condition (F3) is not satisfied, e.g. (1.10) with $p>q=2$. Then one needs to consider a monotonicity-type condition as in [5, 13, 14] given as follows:
  • (F4)
[…]
    Moreover, $\langle f(u),u\rangle\geq 2F(u)$ for every $u\in\mathbb{R}^{2}$.
If $F$ satisfies (F1), (F2), (F4) and $F(u)\gtrsim\Phi(|u|)$ for every $u\in\mathbb{R}^{2}$, then one can work with the Nehari–Pankov manifold for $J$, or Nehari manifold for $\widetilde{J}$:
${\mathcal{N}}:=\{u\in({\mathcal{V}}\oplus{\mathcal{W}})\setminus{\mathcal{W}}:J^{\prime}(u)|_{\mathbb{R}u\oplus{\mathcal{W}}}=0\}=\{u\in\mathcal{V}\setminus\{0\}:\widetilde{J}^{\prime}(u)(u)=0\},$
and build a critical point theory on this constraint as in [5, 6]. We leave the details to the reader.
4. Non-TE modes
In this section, we will work with the space
$X_{{\mathcal{S}}{\mathcal{O}}(2)}:=\left\{u\in X:u=g\star u\text{ for all }g\in{\mathcal{S}}{\mathcal{O}}(2)\right\}.$
For any ${\mathcal{S}}{\mathcal{O}}(2)$-equivariant profile $U\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$, there is a unique decomposition
$U=U_{\rho}+U_{\tau}+U_{\zeta},$
such that $U_{\rho}$, $U_{\tau}$, $U_{\zeta}$ are vector fields given by
$U_{\rho}(x)=\frac{\alpha_{\rho}(x)}{|x|}\begin{pmatrix}x_{1}\\ x_{2}\\ 0\end{pmatrix},\quad U_{\tau}(x)=\frac{\alpha_{\tau}(x)}{|x|}\begin{pmatrix}-x_{2}\\ x_{1}\\ 0\end{pmatrix},\quad U_{\zeta}(x)=\alpha_{\zeta}(x)\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$,
for some ${\mathcal{S}}{\mathcal{O}}(2)$-invariant functions $\alpha_{\rho},\alpha_{\tau},\alpha_{\zeta}\colon\mathbb{R}^{2}\to\mathbb{R}$, cf. [3, Lemma 1]. In fact, $\alpha_{\rho}(x)=\bigl(U_{1}(x)x_{1}+U_{2}(x)x_{2}\bigr)|x|^{-1}$, $\alpha_{\tau}(x)=\bigl(-U_{1}(x)x_{2}+U_{2}(x)x_{1}\bigr)|x|^{-1}$, $\alpha_{\zeta}(x)=U_{3}(x)$ for a.e. $x\in\mathbb{R}^{2}$.
Now for any $u=\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in X_{{\mathcal{S}}{\mathcal{O}}(2)}$ we define
$u_{\rho}(x)=\begin{pmatrix}U_{\rho}\\ \widetilde{U}\end{pmatrix},\quad u_{\tau}(x)=\begin{pmatrix}U_{\tau}\\ \mathbf{0}\end{pmatrix},\quad u_{\zeta}(x)=\begin{pmatrix}U_{\zeta}\\ \mathbf{0}\end{pmatrix}$,
where $\mathbf{0}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}$.
Lemma 4.1.
For every $u\in X_{{\mathcal{S}}{\mathcal{O}}(2)}$ we have $u_{\tau},u_{\rho}+u_{\zeta}\in X_{{\mathcal{S}}{\mathcal{O}}(2)}$ and $u=u_{\rho}+u_{\tau}+u_{\zeta}$, and, for a.e. $x\in\mathbb{R}^{2}$, $u_{\rho}(x)$, $u_{\tau}(x)$, and $u_{\zeta}(x)$ are mutually orthogonal in $\mathbb{R}^{6}$. Moreover, if $u\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, then $u_{\tau},u_{\rho}+u_{\zeta}\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$. If $u\in{\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, then $u_{\tau}=0$ and $u_{\rho}+u_{\zeta}\in{\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$. Finally,
(4.1) $\langle\overset{\circ}{\nabla}\times u_{\rho}(x),\overset{\circ}{\nabla}\times u_{\tau}(x)\rangle=\langle\overset{\circ}{\nabla}\times u_{\tau}(x),\overset{\circ}{\nabla}\times u_{\zeta}(x)\rangle=0$
for a.e. $x\in\mathbb{R}^{2}$.
Proof.
Since $U$ and $\widetilde{U}$ are ${\mathcal{S}}{\mathcal{O}}$-equivariant, $\alpha_{\rho}$, $\alpha_{\tau}$, and $\alpha_{\zeta}$ are ${\mathcal{S}}{\mathcal{O}}$-invariant, hence radial. Since $u_{\rho}(x)$, $u_{\tau}(x)$, and $u_{\zeta}(x)$ are mutually orthogonal in $\mathbb{R}^{6}$, then $u_{\rho},u_{\tau},u_{\zeta}\in L^{\Phi}_{{\mathcal{S}}{\mathcal{O}}(2)}$. Observe that
$u_{\rho}(x)=\begin{pmatrix}\alpha(x)x_{1}\\ \alpha(x)x_{2}\\ 0\\ \widetilde{\alpha}_{1}(x)\\ \widetilde{\alpha}_{2}(x)\\ \widetilde{\alpha}_{3}(x)\end{pmatrix},\quad u_{\tau}(x)=\begin{pmatrix}-\beta(x)x_{2}\\ \beta(x)x_{1}\\ 0\\ 0\\ 0\\ 0\end{pmatrix},\quad u_{\zeta}(x)=\begin{pmatrix}0\\ 0\\ \gamma(x)\\ 0\\ 0\\ 0\end{pmatrix}$
for some $\widetilde{\alpha}_{1},\widetilde{\alpha}_{2}$, and radial $\alpha,\beta,\gamma,\widetilde{\alpha}_{3}$. Note that
$\overset{\circ}{\nabla}\times u_{\rho}(x)=\begin{pmatrix}-k\widetilde{\alpha}_{2}\\ k\widetilde{\alpha}_{1}\\ \partial_{x_{1}}\alpha x_{2}-\partial_{x_{2}}\alpha x_{1}\\ \partial_{x_{2}}\widetilde{\alpha}_{3}+k\alpha x_{2}\\ -k\alpha x_{1}-\partial_{x_{1}}\widetilde{\alpha}_{3}\\ \partial_{x_{1}}\widetilde{\alpha}_{2}-\partial_{x_{2}}\widetilde{\alpha}_{1}\end{pmatrix}=\begin{pmatrix}-k\widetilde{\alpha}_{2}\\ k\widetilde{\alpha}_{1}\\ 0\\ \partial_{x_{2}}\widetilde{\alpha}_{3}+k\alpha x_{2}\\ -k\alpha x_{1}-\partial_{x_{1}}\widetilde{\alpha}_{3}\\ \partial_{x_{1}}\widetilde{\alpha}_{2}-\partial_{x_{2}}\widetilde{\alpha}_{1}\end{pmatrix},\quad\overset{\circ}{\nabla}\times u_{\tau}(x)=\begin{pmatrix}0\\ 0\\ \partial_{x_{1}}(\beta x_{1})+\partial_{x_{2}}(\beta x_{2})\\ k\beta x_{1}\\ k\beta x_{2}\\ 0\end{pmatrix}$,
and since
$(\partial_{x_{2}}\widetilde{\alpha}_{3}+k\alpha x_{2})x_{1}+(-k\alpha x_{1}-\partial_{x_{1}}\widetilde{\alpha}_{3})x_{2}=\partial_{x_{2}}\widetilde{\alpha}_{3}x_{1}-\partial_{x_{1}}\widetilde{\alpha}_{3}x_{3}=0,$
then
$\langle\overset{\circ}{\nabla}\times u_{\rho}(x),\overset{\circ}{\nabla}\times u_{\tau}(x)\rangle=0.$
Moreover
$\overset{\circ}{\nabla}\times u_{\zeta}(x)=\begin{pmatrix}\partial_{x_{2}}\gamma\\ -\partial_{x_{1}}\gamma\\ 0\\ 0\\ 0\\ 0\end{pmatrix}$
and
$\langle\overset{\circ}{\nabla}\times u_{\tau}(x),\overset{\circ}{\nabla}\times u_{\zeta}(x)\rangle=0.$
Therefore, (4.1) holds and
$|\overset{\circ}{\nabla}\times u|^{2}=|\overset{\circ}{\nabla}\times u_{\tau}|^{2}+|\overset{\circ}{\nabla}\times(u_{\rho}+u_{\zeta})|^{2}.$
Hence, if $u\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, then $u_{\tau},u_{\rho}+u_{\zeta}\in{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$. If $u\in{\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$, then $u_{\tau}=0$ and $u_{\rho}+u_{\zeta}\in{\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}$. In view of the Helmholtz decomposition Theorem 2.4, we see that $u_{\tau}\in X_{{\mathcal{S}}{\mathcal{O}}(2)}$, hence $u_{\rho}+u_{\zeta}=u-u_{\tau}\in X_{{\mathcal{S}}{\mathcal{O}}(2)}$. ∎
We consider the action ${\mathcal{S}}\colon X_{{\mathcal{S}}{\mathcal{O}}(2)}\to X_{{\mathcal{S}}{\mathcal{O}}(2)}$ given by
${\mathcal{S}}u:=u_{\rho}-u_{\tau}+u_{\zeta}$
and note that (4.1) yields
$\int_{\mathbb{R}^{2}}|\overset{\circ}{\nabla}\times({\mathcal{S}}u)|^{2}\,dx=\int_{\mathbb{R}^{2}}|\overset{\circ}{\nabla}\times u_{\tau}|^{2}\,dx+\int_{\mathbb{R}^{2}}|\overset{\circ}{\nabla}\times(u_{\rho}+u_{\zeta})|^{2}\,dx=\int_{\mathbb{R}^{2}}|\overset{\circ}{\nabla}\times u|^{2}\,dx.$
Hence we obtain that $\mathcal{S}$} in an isometry and $J|_{X_{{\mathcal{S}}{\mathcal{O}}(2)}}$ is also invariant under $\mathcal{S}$}. Let
$\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}:=\big\{u\in X_{{\mathcal{S}}{\mathcal{O}}(2)}:{\mathcal{S}}u=u\big\}=\big\{u\in X_{{\mathcal{S}}{\mathcal{O}}(2)}:u=u_{\rho}+u_{\zeta}\big\},$
and in view of the Palais principle of symmetric criticality, critical points of $J|_{(X_{{\mathcal{S}}{\mathcal{O}}(2)})^{\mathcal{S}}}$ are critical points of the free functional $J$.
In a similar way we decompose the profile $\widetilde{U}\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$
$\widetilde{U}=\widetilde{U}_{\rho}+\widetilde{U}_{\tau}+\widetilde{U}_{\zeta},$
where $\widetilde{U}_{\rho}$, $\widetilde{U}_{\tau}$, $\widetilde{U}_{\zeta}$ are vector fields given by
$\widetilde{U}_{\rho}(x)=\frac{\widetilde{\alpha}_{\rho}(x)}{|x|}\begin{pmatrix}x_{1}\\ x_{2}\\ 0\end{pmatrix},\quad\widetilde{U}_{\tau}(x)=\frac{\widetilde{\alpha}_{\tau}(x)}{|x|}\begin{pmatrix}-x_{2}\\ x_{1}\\ 0\end{pmatrix},\quad\widetilde{U}_{\zeta}(x)=\widetilde{\alpha}_{\zeta}(x)\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$,
for some ${\mathcal{S}}{\mathcal{O}}(2)$-invariant functions $\widetilde{\alpha}_{\rho},\widetilde{\alpha}_{\tau},\widetilde{\alpha}_{\zeta}\colon\mathbb{R}^{2}\to\mathbb{R}$. Arguing as in Lemma 4.1 we get
$\langle\overset{\circ}{\nabla}\times\widetilde{u}_{\rho}(x),\overset{\circ}{\nabla}\times\widetilde{u}_{\tau}(x)\rangle=\langle\overset{\circ}{\nabla}\times\widetilde{u}_{\tau}(x),\overset{\circ}{\nabla}\times\widetilde{u}_{\zeta}(x)\rangle=0$
for a.e. $x\in\mathbb{R}^{2}$, where
$\widetilde{u}_{\rho}(x)=\begin{pmatrix}U\\ \widetilde{U}_{\rho}\end{pmatrix},\quad\widetilde{u}_{\tau}(x)=\begin{pmatrix}\mathbf{0}\\ \widetilde{U}_{\tau}\end{pmatrix},\quad\widetilde{u}_{\zeta}(x)=\begin{pmatrix}\mathbf{0}\\ \widetilde{U}_{\zeta}\end{pmatrix}$.
Therefore we may define the action $\widetilde{\mathcal{S}}\colon\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\to\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}$ given by
$\widetilde{\mathcal{S}}u:=\widetilde{u}_{\rho}-\widetilde{u}_{\tau}+\widetilde{u}_{\zeta}$
and observe that
$\Big(\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}=\Big\{u=\begin{pmatrix}U\\ \widetilde{U}\end{pmatrix}\in X_{{\mathcal{S}}{\mathcal{O}}(2)}:U=U_{\rho}+U_{\zeta},\;\widetilde{U}=\widetilde{U}_{\rho}+\widetilde{U}_{\zeta}\Big\}.$
Proof of Theorem 1.1. In virtue of Lemma 4.1, we consider the decomposition
$\Big(\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}=\Big(\big({\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}\oplus\Big(\big({\mathcal{W}}_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}$.
Since $\Big(\big({\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}$ is an infinite-dimensional separable Hilbert space, we can consider an orthonormal basis $(v_{n})_{n=1}^{\infty}\subset\Big(\big({\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{\mathcal{S}}\Big)^{\widetilde{\mathcal{S}}}$. Next, for every $n\geq 1$, we define
$Y_{n}:=\bigoplus_{j=1}^{n}\mathrm{span}\{v_{j}\}\quad\text{and}\quad Z_{n}:=\overline{\bigoplus_{j=n}^{\infty}\mathrm{span}\{v_{j}\}}.$
In view of Lemma 3.7, we can directly apply the Fountain Theorem with a Cerami condition [12, Theorem 2.9], cf. [4], and conclude by the Palais principle of symmetric criticality. In particular, we need to prove the existence of two sequences $(r_{n})$ and $(\rho_{n})$ such that $\rho_{n}>r_{n}>0$ for every $n\geq 1$ and the following properties hold:
  • (i)
    $\lim_{n}\left(\inf\left\{\widetilde{J}(v):v\in Z_{n}\text{ and }\|v_{n}\|=r_{n}\right\}\right)=\infty$,
  • (ii)
    For all $n\geq 1$, $\max\left\{\widetilde{J}(v):v\in Y_{n}\text{ and }\|v\|=\rho_{n}\right\}\leq 0$.
We check (i) first. Fix $\varepsilon\in\left(0,\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V\right)$ and let $C_{\varepsilon}>0$ such that (3.1) holds. Then, denoting $c_{\varepsilon}:=\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V-\varepsilon>0$, from (3.2), for every $v\in{\mathcal{V}}$ there holds
$\widetilde{J}(v)\geq c_{\varepsilon}\|v\|^{2}-C_{\varepsilon}\int_{\mathbb{R}^{2}}\Phi(|v|)\,dx.$
Let $\delta\in(0,c_{\varepsilon}/C_{\varepsilon})$. From (N2), there exists $r>0$ such that $\Phi(t)\leq\delta t^{2}$ for all $|t|\leq r$. Now, let $\kappa>1$ be the constant appearing in the characterization of the $\Delta_{2}$ condition, that is, $t\Phi^{\prime}(t)\leq\kappa\Phi(t)$ for all $t\in\mathbb{R}$, cf. Lemma 2.1. Then, there exists $K>0$ such that $\Phi(t)\leq K|t|^{\kappa}$ for all $|t|\geq r$. Observe that this and (N3) yield $\kappa>2$. There holds
$\begin{split}\widetilde{J}(v)&\geq c_{\varepsilon}\|v\|^{2}-C_{\varepsilon}\left(\int_{\{|v|<r\}}\Phi(|v|)\,dx+\int_{\{|v|\geq r\}}\Phi(|v|)\,dx\right)\\ &\geq c_{\varepsilon}\|v\|^{2}-\delta C_{\varepsilon}|v|_{2}^{2}-KC_{\varepsilon}|v|_{\kappa}^{\kappa}\geq(c_{\varepsilon}-\delta C_{\varepsilon})\|v\|^{2}-KC_{\varepsilon}|v|_{\kappa}^{\kappa}.\end{split}$
Letting
$\beta_{n}:=\sup\left\{|v|_{\kappa}:v\in Z_{n}\text{ and }\|v\|=1\right\}$
we obtain, for all $v\in Z_{n}$,
$\widetilde{J}(v)\geq(c_{\varepsilon}-\delta C_{\varepsilon})\|v\|^{2}-KC_{\varepsilon}\beta_{n}^{\kappa}\|v\|^{\kappa}.$
Consequently, choosing
$r_{n}:=\left(\frac{c_{\varepsilon}-\delta C_{\varepsilon}}{2KC_{\varepsilon}\beta_{n}^{\kappa}}\right)^{1/(\kappa-2)}$
we obtain, for all $v\in Z_{n}$ such that $\|v\|=1$,
$\widetilde{J}(v)\geq\frac{c_{\varepsilon}-\delta C_{\varepsilon}}{2}r_{n}^{2}.$
Arguing as in [27, Lemma 3.8] and using the compact embedding $H^{1}_{{\mathcal{S}}{\mathcal{O}}(2)}(\mathbb{R}^{2})^{6}\hookrightarrow\hookrightarrow L^{\kappa}(\mathbb{R}^{2})^{6}$, where
$H^{1}_{{\mathcal{S}}{\mathcal{O}}(2)}(\mathbb{R}^{2})^{6}:=\left\{u\in H^{1}(\mathbb{R}^{2})^{6}:g\star u=u\text{ for all }g\in{\mathcal{S}}{\mathcal{O}}(2)\right\},$
we get that $\lim_{n}r_{n}=\infty$.
Finally, (ii) is a consequence of Lemma 3.4. $\Box$
Proof of (1.12). Let $u$ be a critical point of $J$ and $E$ is of the form (1.1). Observe that
$\langle E,D\rangle$ $=$ $-\varepsilon(x)\omega^{2}|E|^{2}+\chi(\langle|E|^{2}\rangle)|E|^{2}$
$=$ $\big(-V(x)+\chi(\frac{1}{2}|u|^{2})\big)\big(|U|^{2}\cos^{2}(kx_{3}+\omega t)+|\widetilde{U}|^{2}\sin^{2}(kx_{3}+\omega t)\big),$
$\leq$ $\big(-V(x)+\chi(\frac{1}{2}|u|^{2})\big)|u|^{2},$
$\langle B,H\rangle$ $=$ $|B|^{2}=\frac{1}{\omega^{2}}\big|\nabla\times\big(U(x)\sin(kx_{3}+\omega t)-\widetilde{U}(x)\cos(kx_{3}+\omega t)\big)\big|^{2}$
$\leq$ $\frac{1}{\omega^{2}}|\overset{\circ}{\nabla}\times u|^{2}.$
Therefore
${\mathcal{L}}(t)=\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{a}^{a+1}\langle E,D\rangle+\langle B,H\rangle\,dx_{3}\,d(x_{1},x_{2})\leq\frac{1}{2}\bigl(1+\frac{1}{\omega^{2}}\bigr)|\overset{\circ}{\nabla}\times u|^{2}<\infty,$
since $J(u)<\infty$ and $J^{\prime}(u)(u)=0$. $\Box$
Remark 4.2.
One may consider also the action ${\mathcal{S}}\colon X_{{\mathcal{S}}{\mathcal{O}}(2)}\to X_{{\mathcal{S}}{\mathcal{O}}(2)}$ given by
${\mathcal{S}}u:=-u_{\rho}+u_{\tau}-u_{\zeta}$
and note that $\mathcal{S}$} in an isometry and $J|_{X_{{\mathcal{S}}{\mathcal{O}}(2)}}$ is also invariant under $\mathcal{S}$}. Then
$\big(X_{{\mathcal{S}}{\mathcal{O}}(2)}\big)^{(-{\mathcal{S}})}:=\big\{u\in X_{{\mathcal{S}}{\mathcal{O}}(2)}:-{\mathcal{S}}u=u\big\}=\big\{u\in X_{{\mathcal{S}}{\mathcal{O}}(2)}:u=u_{\tau}\big\}$
and in view of the Palais principle of symmetric criticality, critical points of $J|_{(X_{{\mathcal{S}}{\mathcal{O}}(2)})^{(-{\mathcal{S}})}}$ are critical points of the free functional $J$. Arguing as in Section 3 we can find infinitely many solutions of the form $u_{n}=\begin{pmatrix}U_{n}\\ \widetilde{U}_{n}\end{pmatrix}$ with
$U_{n}=\frac{\beta_{n}(x)}{|x|}\begin{pmatrix}-x_{2}\\ x_{1}\\ 0\end{pmatrix},\quad\widetilde{U}_{n}=0$
for some radial $\beta_{n}:\mathbb{R}^{2}\to\mathbb{R}$. Hence $E=U_{n}\cos(kx_{3}+\omega t)$ is a TE-mode. It is not clear however, if for the Kerr nonlinearity $\beta_{n}$ coincide with the solutions obtained in [11] by means of ODE methods.
Acknowledgments. We thank the anonymous reviewer their valuable comments, which helped us improve our manuscript.
Conflict of interest. All authors declare that they have no conflicts of interest.
Ethics approval. Ethical approval was not sought for the present study.
Funding. The authors were partly supported by the National Science Centre, Poland (Grant No. 2020/37/B/ST1/02742). J.S. is a member of GNAMPA (INdAM) and is supported by the GNAMPA project Metodi variazionali e topologici per alcune equazioni di Schrödinger nonlineari.
Data availability statement. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.